What is the value of the reciprocal of (2/3 + 3/2)=
Answer by Mathtut(3670)

2/3+3/2=4/6+9/6=13/6
:
the reciprical of 13/6 is 6/13
:
a number is a reciprical of another number if their product is 1
: